# Hopf orbits and the first ECH capacity

### v2

@@ 2.3. Pseudoholomorphic curves in symplectizations
Let $J$ be a compatible complex structure on the symplectic vector bundle $(\xi,d\lambda)$. As in [93], consider an almost complex structure $\widetilde{J}$ defined on ×$\mathbb{R}\times Y$ by
@@
where $R_{\lambda}$ and $\xi$ are seen as $\mathbb{R}$-invariant objects in ×$\mathbb{R}\times Y$. Then $\widetilde{J}$ is $\mathbb{R}$-invariant. We say that an almost complex structure $\widetilde{J}$ on ×$\mathbb{R}\times Y$ constructed this way is $\lambda$-compatible.
@@
$\widetilde{u}=(a,u):(S\setminus\Gamma,j)\longrightarrow(\mathbb{R}\times Y,\widetilde{J})$
@@
where the supremum is taken over the set of $\phi:\mathbb{R}\to[0,1]$ satisfying $\phi^{\prime}\geq 0$. If the Hofer energy $E(\widetilde{u})<\infty$, then we say that $\widetilde{u}$ is a finite energy curve.
A puncture $z\in\Gamma$ is positive or negative if $a(w)\to+\infty$ or $a(w)\to-\infty$ when $w\to z$, respectively. The puncture $z$ is removable if lim $\limsup|a(w)|<\infty$ when $w\to z$. It turns out that for a finite energy curve, every puncture is positive, negative or removable, and $\widetilde{u}$ can be smoothly extended across a removable puncture; see [93]. So we can assume without loss of generality that $\Gamma$ does not contain removable punctures, and we denote by $\Gamma_{+}$ and $\Gamma_{-}$ the sets of positive and negative punctures, respectively.
@@
Suppose that $\lambda$ is nondegenerate up to action $C$ and that $z$ is a nonremovable puncture of a finite energy curve $\widetilde{u}=(a,u)$ in (×$(\mathbb{R}\times Y,\widetilde{J})$ with Hofer energy $E(\widetilde{u})\leq C$. Let $(s,t)$ be positive holomorphic polar coordinates near $z$. There exist a Reeb orbit $\gamma$ with period $T$ and $d\in\mathbb{R}$ such that $u(s,t)\to\gamma(\epsilon Tt+d)$ in $C^{\infty}(\mathbb{R}/\mathbb{Z},Y)$ as $s\to+\infty$, where $\epsilon=\pm 1$ is the sign of the puncture.
@@
Let $\gamma$ be a Reeb orbit of $\lambda$ with period $T$ and convering multiplicity $k$. A Martinet tube for $\gamma$ is a smooth diffeomorphism $\Psi:\mathcal{N}\to\mathbb{R}/\mathbb{Z}\times\mathbb{D}$ defined on a smooth compact neighborhood $\mathcal{N}$ of $\gamma(\mathbb{R})$ such that:
@@
Suppose that $\lambda$ is nondegenerate up to action $C>0$ and that $z$ is a nonremovable puncture of sign $\epsilon=\pm 1$ of a finite energy curve $\widetilde{u}=(a,u)$ with Hofer energy $E(\widetilde{u})\leq C$. Let $(s,t)$ be positive or negative holomorphic polar coordinates at $z$ when $\epsilon=+1$ or $\epsilon=-1$, respectively. Let $\Psi:\mathcal{N}\to\mathbb{R}/\mathbb{Z}\times\mathbb{D}$ be a Martinet tube for the asymptotic limit $\gamma$ of $\widetilde{u}$ at $z$, and let $s_{0}\gg 1$ such that $|s|\geq s_{0}\Rightarrow u(s,t)\in\mathcal{N}$. Write $\Psi(u(s,t))=(\vartheta(s,t),z(s,t))$ for $|s|\geq s_{0}$. By applying a rotation we can assume $u(s,0)\to x(0)$ as $\epsilon s\to+\infty$.
@@
  • There exist $c,d\in\mathbb{R}$ and a lift $\tilde{\vartheta}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ of $\vartheta(s,t)$ such that
    $\lim_{\epsilon s\to+\infty}\sup_{t\in\mathbb{R}/\mathbb{Z}}e^{r\epsilon s}\left(|D^{\beta}[a(s,t)-Ts-c]|+|D^{\beta}[\tilde{\vartheta}(s,t)-kt-d]|\right)=0$
@@
The alternative $z(s,t)\equiv 0$ can be expressed independently of coordinates as saying that the end of the domain of $\widetilde{u}$ corresponding to the puncture is mapped into the “trivial cylinder” ×$\mathbb{R}\times\gamma$ where $\gamma$ is the asymptotic limit. In this case we say that $\widetilde{u}$ has trivial asymptotic behaviour at the puncture. Otherwise, the asymptotic behaviour is said to be nontrivial at the puncture.
We now recall some topogical invariants introduced in [props2]. Let $\widetilde{u}=(a,u)$ be a finite energy curve in (×$(\mathbb{R}\times Y,\widetilde{J})$. Assume that $\lambda$ is nondegenerate up to action $E(\widetilde{u})$. Let
@@
denote the projection along the Reeb vector field $R_{\lambda}$. It can be shown using the Carleman similarity principle that if $\pi_{\lambda}\circ du$ does not vanish identically, then its zeros are isolated and count positively. Theorem 2.4 further implies that there are finitely many zeros in this case. So if $\widetilde{u}$ is nontrivial, then the algebraic count of zeros of $\pi_{\lambda}\circ du$ is a well-defined non-negative number, which following [props2] we denote by $\operatorname{wind}_{\pi}(\widetilde{u})$.
@@
Note that $\operatorname{wind}_{\infty}(\widetilde{u})$ does not depend on the choice of trivialization $\tau$ of $u^{*}\xi$. In fact, since $\pi_{\lambda}\circ du$ is a section of the complex line bundle $T^{1,0}S\otimes\xi$, it follows from degree theory that
@@
The covering multiplicity ${\rm cov}(\widetilde{u})$ of the plane $\widetilde{u}$ is the covering multiplicity of its asymptotic limit.
@@ 2.4. Pseudoholomorphic curves in strong symplectic cobordisms
Let $(W,\omega)$ be a compact symplectic $4$-manifold for which there exists a Liouville vector field $V$ near $Y=\partial W$ that is transverse to $Y$. The $1$-form $\alpha=i_{V}\omega$ defined near $Y$ satisfies $d\alpha=\omega$ and restricts to a contact form on $Y$ which we denote by $\lambda$. We denote the associated contact structure by $\xi=\ker\lambda\subset TY$. The boundary splits as $Y=Y^{+}\sqcup Y^{-}$, where $V$ points out along $Y^{+}$ and points in along $Y^{-}$. The orientation of $Y$ as the boundary of $W$ agrees with the contact orientation on $Y^{+}$ and disagrees with it on $Y^{-}$. We call $(W,\omega)$ a strong symplectic cobordism that is convex at $(Y^{+},\lambda)$ and concave at $(Y^{-},\lambda)$. The local flow $\varphi^{t}_{V}$ of $V$ defines, for $\varepsilon>0$ small enough, diffeomorphisms
@@
onto neighborhoods $\mathcal{U}^{\pm}$ of $Y^{\pm}$ in $W$. We have $(\Phi^{\pm})^{*}\alpha=e^{a}\lambda$, where here we write $\lambda$ for the pull-back of $\lambda$ under the projection ×$\mathbb{R}\times Y\to Y$. The symplectic form pulls back to
@@
Fix a $\lambda$-compatible almost complex structure $\widetilde{J}$ on ×$\mathbb{R}\times Y$. Consider an almost complex structure $\bar{J}$ on $\overline{W}$ that is $\omega$-compatible on $W$, and that agrees with $\widetilde{J}$ on $[0,+\infty)\times Y^{+}\cup(-\infty,0]\times Y^{-}$ as in (2.10). It follows that $\bar{J}$ is $\overline{\omega}$-compatible. As in [93, sft_comp] we consider a closed Riemann surface $(S,j)$, a finite set $\Gamma\subset S$ and a holomorphic map
@@
We assume the finite energy condition $0<E(\widetilde{u})<\infty$, where the energy $E(\widetilde{u})$ is defined by
@@
Here the supremum is taken over the set of smooth $\phi:\mathbb{R}\to[0,1]$ satisfying $\phi^{\prime}\geq 0$. The finite energy condition implies that punctures are either removable or behave like the nonremovable punctures as explained in §2.3. More precisely, if $(s,t)$ are positive holomorphic polar coordinates at a positive puncture $z\in\Gamma$ then $\widetilde{u}(s,t)\in[0,+\infty)\times Y^{+}$ when $s\gg 1$, and if one writes $\widetilde{u}=(a,u)$ for $s\gg 1$ then $a(s,t)\to+\infty$ as $s\to+\infty$. There is a similar behaviour at negative punctures. Moreover, if $\lambda$ is nondegenerate up to action $E(\widetilde{u})$, then there is an asymptotic limit $P$ at a nonremovable puncture, and all the conclusions of Theorem 2.4 hold. In particular, we can talk about nontrivial versus trivial asymptotic behaviour and asymptotic eigenvalues, see Remark 2.5.
@@
where $\partial_{a},R_{\lambda}\right>$ denotes the subbundle of $T([0,+\infty)\times Y^{+})$ spanned by $\{\partial_{a},R_{\lambda}\}$. There is a $d\lambda$-symplectic frame $\{e_{1},e_{2}\}$ of $(u|_{\mathbb{C}\setminus B_{R}(0)})^{*}\xi$ such that $\{\partial_{a},R_{\lambda},e_{1},e_{2}\}$ extends to a $\overline{\omega}$-symplectic frame of $\widetilde{u}^{*}T\overline{W}$. This frame is unique up to homotopy. If $\lambda$ nondegenerate up to action $E(\widetilde{u})$, then $\widetilde{u}$ has an asymptotic limit $\gamma=(x,T)\in\mathcal{P}(\lambda)$ in $Y^{+}$, and it follows from the asymptotic behaviour that, up to homotopy, $\{e_{1},e_{2}\}$ induces a $d\lambda$-symplectic trivialization $\tau_{\widetilde{u}}$ of $x(T\cdot)^{*}\xi$. Define
(2.16) ${\rm CZ}(\widetilde{u})={\rm CZ}_{\tau_{\widetilde{u}}}(\gamma),$
If the asymptotic behaviour of $\widetilde{u}$ is nontrivial, define
@@
where $\nu$ is the asymptotic eigenvalue of $\widetilde{u}$ at the puncture $\infty$. If the asymptotic behaviour is trivial then we define
(2.18) $\operatorname{wind}_{\infty}(\widetilde{u})=-\infty.$
Under the above nondegeneracy assumption, we define a plane with a positive puncture to be fast if $\operatorname{wind}_{\infty}(\widetilde{u})\leq 1$. Unlike the case of symplectizations, here a fast plane need not be immersed or somewhere injective; the proof of Lemma 2.8 does not carry over because $\operatorname{wind}_{\pi}$ does not make sense in this context. As before, we define ${\rm cov}(\widetilde{u})$ to be the covering multiplicity of the asymptotic limit. If ${\rm cov}(\widetilde{u})=1$ then $\widetilde{u}$ is somewhere injective.
@@ 2.5. The main compactness argument
We now prove Lemma 2.1. We proceed in five steps.
@@
with the symplectic form $\omega$ given by the restriction of the standard symplectic form $\omega_{0}$ on $\Omega_{2}$. We have
$\partial W=Y^{+}-Y^{-},$
where $Y^{+}=\partial\Omega_{2}$ and $Y^{-}=\varphi(\partial\Omega_{1})$. The radial vector field on $\Omega_{2}$ is Liouville, transverse and outward pointing at $Y^{+}$. The pushforward of the radial vector field on $\Omega_{1}$ by $\varphi$ extends to a Liouville vector field defined on a neighbourhood of $Y^{-}$ in $W$, where it is transverse to $Y^{-}$ and inward pointing. Hence we can define the symplectic completion $(\overline{W},\overline{\omega})$ of $(W,\omega)$ as explained in subsection 2.4. As usual, let $\lambda$ denote the contact form on $Y^{+}\cup Y^{-}$. By assumption, $\lambda$ is, up to action $C$, dynamically convex and nondegenerate. Write $\xi=\ker\lambda$. Fix a global $d\lambda$-symplectic trivialization $\tau$ of $\xi$.
Choose any complex structure $J$ on $\xi$ which is $d\lambda$-compatible, and let $\widetilde{J}$ be the corresponding $\lambda$-compatible almost complex structure on ×$\mathbb{R}\times Y$ as in equation (2.10). Let $\bar{J}$ be an almost complex structure on $\overline{W}$ which is $\omega$-compatible on $W$ and agrees with $\widetilde{J}$ on $(-\infty,0]\times Y^{-}\cup[0,+\infty)\times Y^{+}$. In particular, $\bar{J}$ is $\overline{\omega}$-compatible. We need to choose the extension $\bar{J}$ generically, in a way to be specified in Step 4 below.
@@
Let $\gamma_{+}\in\mathcal{P}(Y^{+})$ be a simple Reeb orbit with period $T\leq C$. Let $\mathcal{M}^{\rm fast}(\gamma_{+},\bar{J})$ denote the moduli space of equivalence classes of fast and embedded finite energy planes $\widetilde{u}:(\mathbb{C},i)\to(\overline{W},\bar{J})$ asymptotic to $\gamma_{+}$. Here two planes $\widetilde{u}_{0},\widetilde{u}_{1}$ are declared equivalent if there exist $(A,B)\in\mathbb{C}^{*}\times\mathbb{C}$ such that $\widetilde{u}_{1}(Az+B)=\widetilde{u}_{0}(z)$ for all $z\in\mathbb{C}$. We now show that $\mathcal{M}^{\rm fast}(\gamma_{+},\bar{J})$ is naturally endowed with the structure of a two-dimensional smooth manifold.

### v3

@@ 2.3. Pseudoholomorphic curves in symplectizations
Let $J$ be a compatible complex structure on the symplectic vector bundle $(\xi,d\lambda)$. As in [93], consider an almost complex structure $\widetilde{J}$ defined on $\mathbb{R}\times Y$ by
@@
where $R_{\lambda}$ and $\xi$ are seen as $\mathbb{R}$-invariant objects in $\mathbb{R}\times Y$. Then $\widetilde{J}$ is $\mathbb{R}$-invariant. We say that an almost complex structure $\widetilde{J}$ on $\mathbb{R}\times Y$ constructed this way is $\lambda$-compatible.
@@
[…]
@@
[…]
A puncture $z\in\Gamma$ is positive or negative if $a(w)\to+\infty$ or $a(w)\to$-$\infty$ when $w\to z$, respectively. The puncture $z$ is removable if $\limsup|a(w)|<\infty$ when $w\to z$. It turns out that for a finite energy curve, every puncture is positive, negative or removable, and $\widetilde{u}$ can be smoothly extended across a removable puncture; see [93]. So we can assume without loss of generality that $\Gamma$ does not contain removable punctures, and we denote by $\Gamma_{+}$ and $\Gamma_{-}$ the sets of positive and negative punctures, respectively.
@@
[…]
@@
Let $\gamma$ be a Reeb orbit of $\lambda$ with period $T$ and converging multiplicity $k$. A Martinet tube for $\gamma$ is a smooth diffeomorphism $\Psi:\mathcal{N}\to\mathbb{R}/\mathbb{Z}\times\mathbb{D}$ defined on a smooth compact neighborhood $\mathcal{N}$ of $\gamma(\mathbb{R})$ such that:
@@
For the existence of Martinet tubes the reader is referred to [props1, Lemma 2.3].
@@
Suppose that $\lambda$ is nondegenerate up to action $C>0$ and that $z$ is a nonremovable puncture of sign $\epsilon=\pm 1$ of a finite energy curve $\widetilde{u}=(a,u)$ with Hofer energy $E(\widetilde{u})\leq C$. Let $(s,t)$ be positive or negative holomorphic polar coordinates at $z$ when $\epsilon=+1$ or $\epsilon=-1$, respectively. Let $\Psi:\mathcal{N}\to\mathbb{R}/\mathbb{Z}\times\mathbb{D}$ be a Martinet tube for the asymptotic limit $\gamma$ of $\widetilde{u}$ at $z$, and let $s_{0}\gg 1$ such that $|s|\geq s_{0}\Rightarrow u(s,t)\in\mathcal{N}$. Write $\Psi(u(s,t))=(\vartheta(s,t),z(s,t))$ for $|s|\geq s_{0}$. By applying a rotation we can assume $u(s,0)\to\gamma(0)$ as $\epsilon s\to+\infty$.
@@
  • There exist $c,d\in\mathbb{R}$ and a lift $\tilde{\vartheta}$$\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ of $\vartheta(s,t)$ such that
    $\lim_{\epsilon s\to+\infty}\sup_{t\in\mathbb{R}/\mathbb{Z}}e^{r\epsilon s}\left(|D^{\beta}[a(s,t)-Ts-c]|+|D^{\beta}[\tilde{\vartheta}(s,t)-kt]|\right)=0$
@@
The alternative $z(s,t)\equiv 0$ can be expressed independently of coordinates as saying that the end of the domain of $\widetilde{u}$ corresponding to the puncture is mapped into the “trivial cylinder” $\mathbb{R}\times\gamma$ where $\gamma$ is the asymptotic limit. In this case we say that $\widetilde{u}$ has trivial asymptotic behavior at the puncture. Otherwise, the asymptotic behavior is said to be nontrivial at the puncture.
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 2.4. Pseudoholomorphic curves in strong symplectic cobordisms
Consider a compact symplectic $4$-manifold $(W,\omega)$ for which there exists a Liouville vector field $V$ near $Y=\partial W$ transverse to $Y$. The $1$-form $\alpha=i_{V}\omega$ defined near $Y$ satisfies $d\alpha=\omega$ and restricts to a contact form on $Y$ which we denote by $\lambda$. We denote the associated contact structure by $\xi=\ker\lambda\subset TY$. The boundary splits as $Y=Y^{+}\sqcup Y^{-}$, where $V$ points out along $Y^{+}$ and points in along $Y^{-}$. The orientation of $Y$ as the boundary of $W$ agrees with the contact orientation on $Y^{+}$ and disagrees with it on $Y^{-}$. We call $(W,\omega)$ a strong symplectic cobordism that is convex at $(Y^{+},\lambda)$ and concave at $(Y^{-},\lambda)$. The local flow $\varphi^{t}_{V}$ of $V$ defines, for $\varepsilon>0$ small enough, diffeomorphisms
@@
onto neighborhoods $\mathcal{U}^{\pm}$ of $Y^{\pm}$ in $W$. We have $(\Phi^{\pm})^{*}\alpha=e^{a}\lambda$, where here we still write $\lambda$ for the pull-back of $\lambda$ under the projection $\mathbb{R}\times Y\to Y$. The symplectic form pulls back to
@@
Fix a $\lambda$-compatible almost complex structure $\widetilde{J}$ on $\mathbb{R}\times Y$. Consider an almost complex structure $\bar{J}$ on $\overline{W}$ that is $\omega$-compatible on $W$, and that agrees with $\widetilde{J}$ on $0,+\infty)\times Y^{+}\cup(-\infty,0]\times Y^{-}$ as in (2.10). It follows that $\bar{J}$ is $\overline{\omega}$-compatible. As in [93, sft_comp] we consider a closed Riemann surface $(S,j)$, a finite set $\Gamma\subset S$ and a holomorphic map
@@
[…]
@@
Here the supremum is taken over the set of smooth $\phi:\mathbb{R}\to[0,1]$ satisfying $\phi^{\prime}\geq 0$. The finite energy condition implies that punctures are either removable or behave like the nonremovable punctures as explained in §2.3. More precisely, if $(s,t)$ are positive holomorphic polar coordinates at a positive puncture $z\in\Gamma$ then $\widetilde{u}(s,t)\in[0,+\infty)\times Y^{+}$ when $s\gg 1$, and if one writes $\widetilde{u}=(a,u)$ for $s\gg 1$ then $a(s,t)\to+\infty$ as $s\to+\infty$. There is a similar behavior at negative punctures. Moreover, if $\lambda$ is nondegenerate up to action E(u~)E($\widetilde${u}), then there is an asymptotic limit $P$ at a nonremovable puncture, and all the conclusions of Theorem 2.5 hold. In particular, we can talk about nontrivial versus trivial asymptotic behavior and asymptotic eigenvalues, see Remark 2.6.
@@
where $\left<\partial_{a},R_{\lambda}\right$> denotes the subbundle of $T([0,+\infty)\times Y^{+})$ spanned by $\{\partial_{a},R_{\lambda}\}$. There is a $d\lambda$-symplectic frame $\{e_{1},e_{2}\}$ of $(u|_{\mathbb{C}\setminus B_{R}(0)})^{*}\xi$ such that $\{\partial_{a},R_{\lambda},e_{1},e_{2}\}$ extends to a $\overline{\omega}$-symplectic frame of $\widetilde{u}^{*}T\overline{W}$. Just take, for instance, a frame that extends to a global frame of $u^{*}\xi$. This frame is unique up to homotopy: any two such frames would have relative Maslov index equal to zero along any loop in their domain of definition $\mathbb{C}$. If $\lambda$ nondegenerate up to action E(u~)E($\widetilde${u}), then $\widetilde{u}$ has an asymptotic limit $\gamma$ in $Y^{+}$, of period $T>0$, and it follows from the asymptotic behavior that, up to homotopy, $\{e_{1},e_{2}\}$ induces a $d\lambda$-symplectic trivialization $\tau_{\widetilde{u}}$ of $\gamma(T\cdot)^{*}\xi$. Define
(2.16) $\operatorname{CZ}(\widetilde{u})=\operatorname{CZ}_{\tau_{\widetilde{u}}}(\gamma),$
If the asymptotic behavior of $\widetilde{u}$ is nontrivial, define
@@
where $\nu$ is the asymptotic eigenvalue of $\widetilde{u}$ at the puncture $\infty$. If the asymptotic behavior is trivial then we define
[…]
@@ 2.5. The main compactness argument
We now prove Lemma 2.1. Before diving into the argument, let us sketch the main steps.
Firstly, we use the lemma holds under the assumptions that $\varphi(\Omega_{1})\subset{\rm int}(\Omega_{2})$ and $\Omega_{2}$ is nondegenerate up to action $C$, to show the version of the lemma when these assumptions are dropped. This is Step 0. The remaining argument is to show the lemma when $\varphi(\Omega_{1})\subset{\rm int}(\Omega_{2})$ and $\Omega_{2}$ is nondegenerate up to action $C$. The idea is: (1) consider the symplectic cobordism from $\partial\Omega_{2}$ to $\varphi(\partial\Omega_{1})$ induced by $\Omega_{2}$, which has a convex boundary at $\partial\Omega_{2}$ and a concave boundary at $\varphi(\partial\Omega_{1})$, (2) use existence results for fast planes from [fast, hryn_jsg] to see that some moduli space of fast planes in this cobordism is nonempty, (3) “push” the fast planes in this moduli space “inside” the concave boundary $\varphi(\partial\Omega_{1})$, (4) show that this family of planes SFT-converges to cylindrical building consisting of a cylinder in the cobordism, and a fast plane in the symplectization of $\varphi(\partial\Omega_{1})$ which is asymptotic to the desired Hopf orbit in $\varphi(\partial\Omega_{1})$.
Step 0
First assume that Lemma 2.1 holds under the additional assumptions that $\varphi(\Omega_{1})\subset{\rm int}(\Omega_{2})$ and that $\Omega_{2}$ is nondegenerate up to action $C$. In this case the lemma provides the orbit $\gamma\in\mathcal{P}_{\rm Hopf}(\partial\Omega_{1})$ with a bound $\mathcal{A}(\gamma)<\mathcal{A}_{\rm Hopf}(\Omega_{2})$. Using this, we will now prove the version of Lemma 2.1 where these extra assumptions are dropped.
Consider a sequence $\gamma_{n}\in\mathcal{P}_{\rm Hopf}(\partial\Omega_{2})$ satisfying $\mathcal{A}(\gamma_{n})\to\mathcal{A}_{\rm Hopf}(\Omega_{2})$. Choose $\epsilon_{n}>0$, $\epsilon_{n}\to 0$ an arbitrary sequence. We follow a scheme similar to Step 4 in the proof of Theorem 1.9. Note that $\tilde{\gamma}_{n}=e^{\epsilon_{n}}\gamma_{n}$ is a Hopf orbit in $e^{\epsilon_{n}}\partial\Omega_{2}$. Let $\Omega^{n}_{2}$ be a small $C^{\infty}$-perturbation of $e^{\epsilon_{n}}\Omega_{2}$ such that:
  • $\Omega^{n}_{2}$ is nondegenerate and dynamically convex up to action $C$.
  • $\tilde{\gamma}_{n}\subset\partial\Omega^{n}_{2}$.
  • $\partial\Omega^{n}_{2}\to\partial\Omega_{2}$ in $C^{\infty}$ as hypersurfaces.
  • $\Omega_{2}\subset{\rm int}(\Omega^{n}_{2})$.
In particular, $\tilde{\gamma}_{n}\in\mathcal{P}_{\rm Hopf}(\partial\Omega^{n}_{2})$. Choose $\delta_{n}\in(0,\epsilon_{n})$ such that $\Omega_{2}\subset{\rm int}(e^{-\delta_{n}}\Omega^{n}_{2})$. Note that $e^{-\delta_{n}}\tilde{\gamma}_{n}=e^{\epsilon_{n}-\delta_{n}}\gamma_{n}$ belongs to $\mathcal{P}_{\rm Hopf}(e^{-\delta_{n}}\partial\Omega_{2}^{n})$. By the assumed version of Lemma 2.1, there exist periodic orbits $\hat{\gamma}_{n}\in\mathcal{P}_{\rm Hopf}(\partial\Omega_{1})$ satisfying
$\mathcal{A}(\hat{\gamma}_{n})<\mathcal{A}_{\rm Hopf}(e^{-\delta_{n}}\Omega^{n}_{2})\leq e^{\epsilon_{n}-\delta_{n}}\mathcal{A}(\gamma_{n})\to\mathcal{A}_{\rm Hopf}(\Omega_{2}),\qquad{\rm CZ}(\hat{\gamma}_{n})=3.$
A direct application of Lemma 2.2 to the sequence $\hat{\gamma}_{n}$ in $\Omega_{1}$ establishes the existence of $\gamma\in\mathcal{P}_{\rm Hopf}(\Omega_{1})$ satisfying $\mathcal{A}(\gamma)\leq\mathcal{A}_{\rm Hopf}(\Omega_{2})$ and $\operatorname{CZ}(\gamma)=3$. The proof of Step 0 is complete.
With Step 0 established, we now proceed with the proof of Lemma 2.1 under the additional assumptions that $\Omega_{2}$ is nondegenerate up to action $C>\mathcal{A}_{\rm Hopf}(\Omega_{2})$, and that $\varphi(\Omega_{1})\subset{\rm int}(\Omega_{2})$. In this case we wish to find $\gamma\in\mathcal{P}_{\rm Hopf}(\Omega_{1})$ satisfying $\operatorname{CZ}(\gamma)=3$ and $\mathcal{A}(\gamma)<\mathcal{A}_{\rm Hopf}(\Omega_{2})$.
@@
with symplectic form $\omega=\omega_{0}$. We have
$\partial W=Y^{+}-Y^{-}$
where $Y^{+}=\partial\Omega_{2}$ and $Y^{-}=\varphi(\partial\Omega_{1})$. The radial vector field on $\Omega_{2}$ is Liouville, transverse and outward pointing at $Y^{+}$. The pushforward of the radial vector field on $\Omega_{1}$ by $\varphi$ extends to a Liouville vector field defined on a neighbourhood of $Y^{-}$ in $W$, where it is transverse to $Y^{-}$ and inward pointing. Hence, we can define the symplectic completion $(\overline{W},\overline{\omega})$ of $(W,\omega)$ as in § 2.4. As usual, let $\lambda$ denote the contact form on $Y^{+}\cup Y^{-}$. By assumption, $\lambda$ is, up to action $C$, dynamically convex and nondegenerate. Write $\xi=\ker\lambda$. Fix a global $d\lambda$-symplectic trivialization $\tau$ of $\xi$.
Choose any complex structure $J$ on $\xi$ which is $d\lambda$-compatible, and let $\widetilde{J}$ be the corresponding $\lambda$-compatible almost complex structure on $\mathbb{R}\times Y$ as in equation (2.10). Let $\bar{J}$ be an almost complex structure on $\overline{W}$ which is $\omega$-compatible on $W$ and agrees with $\widetilde{J}$ on $(-\infty,0]\times Y^{-}\cup[0,+\infty)\times Y^{+}$. In particular, $\bar{J}$ is $\overline{\omega}$-compatible. We need to choose the extension $\bar{J}$ generically, in a way to be specified in Step 4 below.
@@
Let $\gamma_{+}$ be a simple $\lambda$-Reeb orbit in $Y^{+}=\partial\Omega_{2}$ with period $T\leq C$. Let $\mathcal{M}^{\rm fast}(\gamma_{+},\bar{J})$ denote the moduli space of equivalence classes of fast and embedded finite energy planes $\widetilde{u}:(\mathbb{C},i)\to(\overline{W},\bar{J})$ asymptotic to $\gamma_{+}$. Two planes $\widetilde{u}_{0},\widetilde{u}_{1}$ are declared equivalent if there exist $(A,B)\in\mathbb{C}^{*}\times\mathbb{C}$ such that $\widetilde{u}_{1}(Az+B)=\widetilde{u}_{0}(z)$ for all $z\in\mathbb{C}$. We now show that $\mathcal{M}^{\rm fast}(\gamma_{+},\bar{J})$ is naturally endowed with the structure of a two-dimensional smooth manifold.